MARS DESERT RESEARCH STATION

Commander Report – February 21st

Dear Earth,


We are currently living our ninth Sol on Mars, and fortunately the
weather was kind enough with us to allow a three-hour EVA.

Louis Mangin, Louis Maller, Victoria and Simon went out at 9 AM and
began by looking for a good place to anchor the balloon (this maneuver
being planned for a later and less windy EVA). The plan is to let the
balloon fly during an entire Sol, at about 80 meters of height (260
ft) and attached to a heavy rock. During the night, it will slowly go
down, and we will pick it up on the next morning to recover 24 hours
of atmospheric data. To conduct that experiment, we need a long period
with less than 10mph of wind, so we keep our fingers crossed for the
next Sols.

The EVA team finally found the right place, in a plain North-East of
the Hab, visible from its windows. After that, they took two ATVs and
a rover to explore and take pictures at White Rock Canyon, which was
discovered during yesterday’s EVA. Meanwhile, Xavier learned by heart
the datasheet of our electric supply, Mouâdh prepared a delicious rice
with tomato sauce (rice is the only food we have in large quantity,
beside the tremendous amount of snacks of all kinds), and I made a new
funnel from cans, for the refilling of the ATVs’ oil.

The wind rose around 11 AM and went on during the whole Sol. It was as
if the Hab could be blown away at any time. Sometimes the noise
covered our usual musical ambiance, but the work continued,
undisturbed by the elements. We had an important debate in the
afternoon, about our rationing of the breakfast, until we found a
brand new jar of active yeast well hidden on the shelves (our previous
jar vanished with yesterday’s bread). We are saved!

My report was almost over when we realized that the sky was finally
Mars-like: the orange-red clouds in the evening sky were spectacular
and gave us an impression of fireworks’ grand finale, after the wind
storm.


Ad Astra!

Arthur Lillo
Commander of the stunned Crew 175

GreenHab Report – February 20th

Green Hab Report  – Sol 8
Report written by: Victoria DA-POIAN (Crew Biologist)

Date : 02/20/2017

Functionality: The heater in the green hab is working well. Today was a nicer day than yesterday! I checked the temperature in the GreenHab this morning. It was around 27 Celsius degrees around 11:00 AM while the temperature in the tunnel was around 6.8 Celsius degrees. I switched on the cooler (on stage 1). The GreenHab temperature was 21 Celsius degrees at 5:30PM while it was 12.6 Celsius degrees in the tunnel. I watered the seedlings this morning and this afternoon too.


Status: The existing seedlings in the greenhab are continuing to grow well. There are spinach, lettuce, radish, and beans growing very well in the small pots.

The Vegidair has been installed 4 days ago and is functionning very well (cf photos). We were able to see more sprouts of lettuce and of « roquette » today in the Vegidair but not in the similar pots I put in the GreenHab.

Sol Summary – February 20th

SOL SUMMARY REPORT SOL 8
SOL: 8
Person filling out Report: Louis MALLER, XO
Summary Title: White Rock Canyon
Mission Status: interesting exploration EVA, all systems go, work on experiments ongoing
Sol Activity Summary: sport, EVA, science work on different experiments
Look Ahead Plan: As this weather should improve in the following next few days, the crew is looking forwards to further testing of the balloon.
Anomalies in work: heavy fogging, maybe due in part to the very humid air
Weather: foggy in the early morning and then improved a lot
Crew Physical Status: Crew feeling well
EVA: EVA #8 checked on the seismometer and explored the White Rock Canyon.

EVA Report – February 20th

EVA Report

EVA#8

Crew members: Mouâdh Bouayad (EVA leader), Arthur Lillo, Louis Maller and Xavier Rixhon (EVA buddies)

Habcom : Simon Bouriat

Departure time:  9:05 AM

Return time: 11:20 AM

Vehicles: 1 rover and 2 ATVs

Location: South of the Hab (12N 518500, 4250000) and white Rock Canyon (12N 520250, 4247500).

Purposes:

  • White Rock Canyon exploration
  • Seismology experiment: verification of the instruments’ proper functioning

Summary:

We did the engineering check, and we filled the black tank to 90%. It took us about 20 minutes to complete the tasks. After that, we took the tank out of the rover Phobos, took Deimos and two ATVs.

We first went to check the seismology experiment, South on the Main Road, and the seismometer had slightly moved; we moved it back the way it should be. Otherwise everything functioned. We then took the rover and ATVs again and headed to White Rock Canyon. As there was fog in the helmets, we had to drive quite slowly.

After 15 minutes, we finally arrived to the White Rock Canyon. We pulled over and went to explore the canyon at 10:05 AM. After 40 minutes of exploration, we figured out that Xavier didn’t have air supply anymore. Therefore, we plugged one of Arthur’s air pipe to Xavier’s helmet, and I decided to go back to the Hab. They thus had to walk together back to the rover. Once there, we put Arthur’s pipe back where it belonged.

We left the White Rock Canyon at 10:45 AM, and arrived at the Hab at 11:05 AM. We then finished the engineering check of the rover and ATVs. It was quite hard to refill the ATVs with fuel. We then got back to the Hab, at 11:20 AM.

Crew Photos – February 20th

Vegidair in action
Simon helping Louis
White rock canyon 2
White rock canyon 1
Reaching White rock canyon
Landscape
Forced pairw 

Journalist Report – February 20th

Journalist report, 02/20/17 – Sol 8: First shower.

Today’s EVA was led by Mouadh, followed by Arthur, Louis, and Xavier. They left at 9:00, refilled the water tank, checked the seismometer, and went to White Rock canyon to explore it. The weather was once again foggy, but it was clearing when they left. The seismometer had moved a bit, because of the rain from the past few days, making the ground softer. Having picked up the USB key full of data, they went back to their rovers and went to the canyon. It was very wet in there. Louis took beautiful photos from above, before Xavier’s backpack stopped blowing air (due to a bad reload as we discovered after). He was almost blind, and more important, suffocating if we were on Mars. That is why, to respect the simulation constraints, they applied a common protocol in diving: Arthur shared one of his pipes with him, on the way back, forcing them to move as Siamese twins. The EVA stopped thirty minutes earlier because of this failure, but the good point was the quick reaction of the team.

This afternoon, we did for the first time another sport session I brought from my rower’s training, to keep the crew fit. Then, I had my first shower after eight days. We managed to save water as much as possible and Xavier and I were the last two ones having had none. We were trusting last crew’s recommendation about it, that was: “take a minimum number of showers”. But now that we are monitoring more precisely our water consumption, the conclusion was clear: it was ridiculous, as a (quick) shower uses less water than a flush… So that one represents around 3% of one day consumption, while flushing represents between 35% and 40%.

Anyway, I now am clean, and already motivated for the next 11 sols that remain! We already have great plans for the following ones, and I am not in a hurry to leave. I will lead tomorrow’s EVA for the first time, which goal will be to find a place to deploy the balloon for a whole day.

 

Louis MANGIN, crew journalist MDRS 175

Commander Report – February 20th

Dear Earth,


Sol 8: we have been here for an entire week now, and this morning we
woke up with the foggiest weather we had since our arrival.

Mouâdh took the commandership of our morning EVA, leading Louis
Maller, Xavier and I to the southernmost part of our map: White Rock
Canyon. The mist was nearly gone when we arrived there, but on the
inside part of the helmet glass, the fog was terrible. We could only
drive at a very low speed to stay on the Main Road. Despite its not so
impressive size, the canyon was stunning. We discovered orange-red
water when walking along the valley, and lots of colored strata with
various orientations.

Then, the plot thickened: deep inside the canyon, Xavier realized that
its backpack had stopped sending him fresh air. Mouâdh reacted quickly
and wisely by taking the decision to recover to the rovers and to head
back to the Hab. Xavier and I agreed on a solution inspired by our
scuba diving experience to provide him with air: I plugged one of my
air-supply tubes to his helmet, keeping the other one on my own
helmet. That way, Xavier could not only breathe fresh air, but also
see through his helmet, the fog being cleaned by the airstream. We
walked side by side on our way back to the rovers, and then the EVA
team came back to the Hab after almost two and a half hours outdoors,
with the sun finally showing up. This experience was quite inspiring,
indeed afterward we defined some emergency protocols regarding this
kind of situation, likely to happen again if we forget to turn off the
backpack when they are recharging.

In the afternoon, we tried a new physical training proposed by Louis
Mangin, inspired by his usual rowing training. Demanding, but we felt
good after this physical effort, and it was an excellent teambuilding
activity. I am glad to see that even after a whole week in here the
crew is still dynamic and motivated! Ahead of us, we still have about
twelve Sols of mission, which is the typical duration of a rotation at
MDRS.


Ad Astra!

Arthur Lillo,
Commander of the sportive Crew 175

EVA Report – February 19th

EVA report:

Crew members: Simon Bouriat (EVA leader), Victoria Da-Poian, Arthur Lillo and Louis Mangin (EVA buddies)

Location:

  • East of the Hab (Marble Ritual, 12N 518700, 4250700)

Departure time:  9:00 AM

Return time: 10:05 PM

Vehicles used: None

Duration: 1:05 hours

Summary:
We first walked from the Hab to Marble Ritual to test the radio relay built yesterday. In fact, we know that the walkie-talkies start to malfunction around this area because of the distance. So we went there and split the group into two pairs. Two were supposed to keep the radio relay with them and contact the Hab and the two others had to keep walking until they lost contact with the hab. Thanks to the radio relay, this last pair should have been able to contact the Hab. This experiment was a failure for different reasons. First of all, it took a long time to reach Marble Ritual as our helmets were covered of fog. Secondly, the walkie-talkies were too sensitive to the terrain. Most of the time, the HabCom was able to contact the second pair (who was walking) and the two pairs weren’t able to talk to each other. We think that these issues have been caused by the air humidity and by interferences from the radio relay. We will work again to fix that and will try it during a sunny EVA. However, our commander succeeded to use his sextant for his positioning experiment during the EVA, even if the temperature tended to decalibrate it. The fog on our helmets started to be a big issue so we finally got back to the Hab as nobody could see and, so, drive the rovers.

Commander Report – February 19th

Dear Earth,

We began Sol 7 with the same bad weather as yesterday.

Today’s EVA was shortened like the last one. The rain from last night
had left a very muddy environment, where it was not safe to take the
rovers. Simon was in charge of the EVA, with Victoria, Louis Mangin
and I. We went East of the Hab to determine the range of the walkie
talkies and test the radio repeater built yesterday. Unfortunately,
the results were absolutely not concluding, probably due to the
weather: even with a line of sight, sometimes we could not communicate
with each other while the Hab was hearing us. The second experiment
was the marine sextant, with which we measured the angles between
distant objects (the Hab, a remarkable hill…), to determine our
position on the map afterward. The significant humidity in the air
covered our helmets with fog before long, preventing us from seeing
where we stepped. Thus, after an hour out of the Hab, Simon took the
decision to abort the EVA and come back to the main airlock.

Unlike the EVA, the debriefing was efficient. There are much more
lessons to learn from an aborted mission than from a successful one.
We took this opportunity to define protocols for foggy situations
(guidance through radio), or non working radio (non-verbal
communication inspired by diving codes).

The afternoon brought a sunnier touch on our activities, allowing us
to watch our first (and unlikely) Martian rainbow. The experiments
went on normally, with interesting results. The sprouts in Victoria’s
Vegidair grow quicker than the ones in the GreenHab; the coordinates
on the map seem to correspond to our position when using the marine
sextant; the Aquapad petri dishes show without a doubt the destruction
of all bacterial life in boiled water compared with tap water (our
teachers did not lie to us!).

Apart from the crew, there is another busy inhabitant in the Hab: no
rest for the bread machine, the “perpetual bread” is on its way. The
psychological effect of eating fresh loaf everyday is undoubtedly
positive, and I am convinced that a shortage of flour would cause the
mutiny of my crew. I keep an eye on the stock, and an escape plan just
in case.


Ad astra!

Arthur Lillo,
Commander of the EVA-aborting Crew 175

GreenHab Report – February 19th

Green Hab Report  – Sol 7
Report written by: Victoria DA-POIAN (Crew Biologist)

Date : 02/19/2017

Functionality: The heater in the green hab is working well. Today was a rainy, cloudy and cold day ! I checked the temperature in the GreenHab this morning. It was around 21 Celsius degrees around 11:00 AM while the temperature in the tunnel was around 6.8 Celsius degrees. I did not switch on the cooler. The GreenHab temperature was 18 Celsius degrees at 5:00PM while it was 6.6 Celsius degrees in the tunnel. I watered the seedlings this morning and this afternoon too.

Status: The existing seedlings in the greenhab are continuing to grow well. There are spinach, lettuce, radish, and beans growing very well in the small pots.

The Vegidair has been installed 3 days ago and is functionning very well. We were able to see more sprouts of lettuce and of « roquette » today in the Vegidair but not in the similar pots I put in the GreenHab.